1. The ABC bank has a deposit base on which it holds excess reserves of $42 Million at a reserve rate which relates to a money multiplier (MM) of 5. On January 2, 2010 Mwen-Mwen, a customer of ABC bank issued a $20 Million check drawn on its account, to Bushrod Manufacturing Company (BMC), which it deposited into its account with Satiah Bank.

 



1
Expert's answer
2021-09-23T17:37:19-0400

Solution:

a.). The reserve ratio:

Money Multiplier = "\frac{1}{Reserve\; Ratio}"


5 = "\frac{1}{R\; R}"


Reserve ratio = "\frac{1}{5} = 20\%"


Reserve ratio = 20%

 

b.). The initial deposit base prior to the $20 Million transactions:


Change in money supply = Initial Excess Reserves "\times" Money Multiplier = 42M x 5 = 210M


The initial deposit base prior to the $20 Million transactions = 210M